The manager of the Postal Services claims that 98 percent of their mail are delivered on time. He wants to test at the 1% significance level to determine whether the true proportion is less than 98 percent. i. Give the null and alternative hypothesis of this test. [2] ii. Determine the critical value(s) of this test. [2] iii. Compute the value of the test statistic. [2] iv. State the decision rule. [1] v. Give your decision based on the available sample evidence. [1] vi. Hence, state your conclusion


1
Expert's answer
2021-11-02T18:22:23-0400

i. The following null and alternative hypotheses for the population proportion needs to be tested:

"H_0:p\geq0.98"

"H_1:p<0.98"

This corresponds to a left-tailed test, for which a z-test for one population proportion will be used.


ii. Based on the information provided, the significance level is "\alpha = 0.01

," and the critical value for a left-tailed test is "z_c = -2.3263."

The rejection region for this left-tailed test is "R = \{z: z < -2.3263\}."


iii. The z-statistic is computed as follows:


"z=\dfrac{\hat{p}-p_0}{\sqrt{\dfrac{p_0(1-p_0)}{n}}}"

iv. If it is observed that "z < z_c = -2.3263," it is then concluded that the null hypothesis is rejected.

Using the P-value approach: The p-value is "p=P(Z<z)," and if "p<0.01=\alpha," it is concluded that the null hypothesis is rejected.


v. Let "N=500, X=480." Then


"\hat{p}=\dfrac{480}{500}=0.96"


"z=\dfrac{\hat{p}-p_0}{\sqrt{\dfrac{p_0(1-p_0)}{n}}}=\dfrac{0.96-0.98}{\sqrt{\dfrac{0.98(1-0.98)}{500}}}""=-3.1944"

Since it is observed that "z =-3.1944 < z_c = -2.3263," it is then concluded that the null hypothesis is rejected.

Using the P-value approach: The p-value is "p=P(Z<-3.1944)=0.0007," and if "p=0.0007<0.01=\alpha," it is concluded that the null hypothesis is rejected.


vi. Therefore, there is enough evidence to claim that the population proportion "p" is less than 0.98, at the "\alpha = 0.01" significance level.